You are building a half pipe in the snow the length will be 60 meters and the width will be 8 meters how much snow will uou have to remvove?
Answer by stanbon(75887)

You are building a half pipe in the snow the length will be 60 meters and the width will be 8 meters how much snow will uou have to remove?
----------
It's (1/2) the vertical surface of a cylinder::
----
(1/2)perimeter = (1/2)2*pi*r = pi*r
----
Ans: surfact of the half-pipe = pi*r*60 = 60*4*pi = 240pi sq meters
-------
Cheers,
Stan H.
-------